OpenStudy (anonymous):

two years ago, Greg was three times as old as his brother. in 4 years, he will be only twice his brother's age. how old is his brother now?

OpenStudy (anonymous):

let "g" be the age of greg let "b" be the age of his brother 2 years ago, their ages would be "g-2" and "b-2" 3(b-2)=g-2 in 4 years, their ages would be "g+4" and "b+4" 2(b+4)=g+4 now you have 2 equations with 2 unknowns, so just solve for b

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you understand where i got the equations from?

OpenStudy (anonymous):

yes, im trying to solve it now.

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

give me a sec, i did not attempt to solve

OpenStudy (anonymous):

3(b-2)=g-2 2(b+4)=g+4 3b-6=g-2 2b+8=g+4 3b-4=g 2b+4=g

OpenStudy (anonymous):

and the answer issssss?